OpenStudy (anonymous):

Complete the square: x2 + 6x + __ How do you solve this?

OpenStudy (anonymous):

x^2

OpenStudy (anonymous):

not x2

OpenStudy (anonymous):

sorry

OpenStudy (anonymous):

u have to divide 6 by 2, then u square tht number to get ur completed square

OpenStudy (anonymous):

so 1.73?

OpenStudy (anonymous):

6/2, 3 sqrt

OpenStudy (anonymous):

6/2 = 3 3^2 = 9 so 9 is the answer

OpenStudy (anonymous):

6 divided by 2 =3 now square tht 3^2=9

OpenStudy (anonymous):

AH. Thank you.

OpenStudy (anonymous):

gotcha. Thanks a lot guys!